OpenStudy (anonymous):

Find the set of values of p such that px^2+2x+4p < 0 for all x.

OpenStudy (anonymous):

if this is true for all x then its D <0

OpenStudy (anonymous):

Sorry, if i could get some working. I would appreciate it.

OpenStudy (anonymous):



OpenStudy (anonymous):

does it make sense?

OpenStudy (anonymous):

y= px^2+2x+4p and y<0 for all x

OpenStudy (anonymous):

Sorry, i have no idea where you are coming from?

OpenStudy (anonymous):

ok...u have written some quadratic expression...px^2+2x+4p...right?

OpenStudy (anonymous):

Its under the chapter of quadratics. And yes i have written that.

OpenStudy (anonymous):

u know that curve of a quadratic is a parabola?

OpenStudy (anonymous):

Assume all medium levels of knowledge.

OpenStudy (anonymous):

ok....

OpenStudy (anonymous):

Im from australia and here i do the highest level of maths for my respective grade, 11.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

how many roots a quadratic equation has?

OpenStudy (anonymous):

ok....let's draw the curve y = x^2 +3x+4

OpenStudy (anonymous):

are you comfortable with this?

OpenStudy (anonymous):

Yes, I can understand that much. I'm the one in the class who always wants to know why. Why do we remove the p for now? Sorry if its a simple question, i just want to be able to repeat this to someone else if they ask.

OpenStudy (anonymous):

i am not saying that it's a simple question...and simplicity depends on whether one knows or not? even if I had not known to solve this then it wud have been a difficult question for me

OpenStudy (anonymous):

let's draw the curve y = x^2 +3x+4

OpenStudy (anonymous):

So we remove the pronumeral which they wish us to find?

OpenStudy (anonymous):

we draw the curve y = x^2 +3x+4

OpenStudy (anonymous):

Do you mean y=x^2+2x+4?

OpenStudy (anonymous):

still there?

OpenStudy (anonymous):

Yes. Im here and willing to listen.

OpenStudy (anonymous):

isolate the terms with 'p'

OpenStudy (anonymous):

factor out p

OpenStudy (anonymous):

divide both sides by the factor of p (it will be negative so the <> sign must reverse)

OpenStudy (anonymous):

p< -2x/(x^2+4)

OpenStudy (anonymous):

Im with you.

OpenStudy (anonymous):

check the range of -2x/(x^2+4) to find allowed values for p

OpenStudy (anonymous):

Range as in Domain and Range?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

In terms of x or p?

OpenStudy (anonymous):

-2x/(x^2+4) has an absolute min

OpenStudy (anonymous):

if you're not doing calc. yet... hmm.

OpenStudy (anonymous):

are you?

OpenStudy (anonymous):

No I am doing calc.

OpenStudy (anonymous):

ah easy then

OpenStudy (anonymous):

find the min of -2x/(x^2+4)

OpenStudy (anonymous):

p is any value less than that min

OpenStudy (anonymous):

Its in a quadratic section though, so I would prefer not to use calc.

OpenStudy (anonymous):

hmm yeah I think you can use the vertex form that someone mentioned briefly before deleting the comment to do it a different way...